### v1

@@ 5.3 Structural Robustness and Task-Scale Sensitivity
We first group benchmark tasks by the size of their ground-truth patches and evaluate model pass@16 within each bin. As shown in Figure 8, most models exhibit sharp performance degradation as the size of the target edit increases. While all models achieve moderate to high success rates on small-scale edits (1–5 lines), performance drops significantly for medium (11–20 lines) and large-scale (>20 lines) tasks. In particular, o3-mini is the only model to retain non-trivial success on edits exceeding 50 lines, with a pass@16 about 25%, while all other models collapse to near-zero in this regime. These results highlight the inability of most LLMs to scale beyond localized modifications.
Figure 8: Model pass@16 success rate grouped by ground-truth patch size. Larger edits cause sharp performance degradation for most models. Only o3-mini maintains partial success on tasks with more than 20 lines of changes.
Patch Length Alignment with Ground-Truth.
Figure 9 evaluates whether model-generated patches exhibit scale-aware behavior by comparing their length to the corresponding human edits. Most models over-edit small tasks (1–5 lines), producing longer-than-needed patches. As task size increases, generated patch sizes tend to align more closely with ground truth—except in models like GPT-4o and Doubao 1.5 Pro, which increasingly under-edit larger patches. This under-generation corresponds to their low semantic success on structurally complex tasks.
In contrast, o3-mini and other strong models (Claude Sonnet, DeepSeek variants) exhibit broader variance and better coverage across patch sizes. Their outputs span both under- and over-editing, reflecting more diverse structural hypotheses. Notably, o3-mini frequently over-edits larger patches, which increases numeric deviation but improves semantic success by avoiding missing components—the dominant failure mode among conservative models. These results suggest that scale diversity, rather than strict length matching, is more critical for robust proof engineering performance.
@@
Stratified Performance by Task Type, Difficulty, and Domain.
To further investigate structural sensitivity, we evaluate model performance across different task categories, difficulty levels, and mathematical domains. Figure 10 presents a unified heatmap showing evaluation pass@16 for each model across these dimensions. Tasks labeled as bug fixes are handled reasonably well by all models, likely due to their localized and low-risk nature. Refactor tasks reveal stronger model differentiation: Claude variants perform relatively well due to their structural preservation bias, whereas GPT-4o and Doubao underperform. Feature addition tasks are the most challenging, requiring new definitions, theorems, and often multi-declaration coherence; only o3-mini exhibits notable success in this category.
Difficulty-based analysis further reveals a sharp capability cliff. While all models succeed on easy tasks, only o3-mini retains significant performance on medium and hard cases. Hard tasks, involving large edits or cross-declaration dependencies, remain virtually unsolved by all models except o3-mini. Finally, we observe domain-based variation: linear algebra and category theory consistently challenge most models, while tasks from algebra or data modules are comparatively tractable. These differences reflect varying structural depth and abstraction levels across formal math subfields.
Figure 10: Heatmap of evaluation pass@16 success rates across different task types, difficulty levels, and mathematical modules. Column widths represent the proportion of each class in the test set. Color intensity indicates model performance.
Summary.
Taken together, these results indicate that most current models lack the structural robustness necessary for realistic proof engineering. Their success is largely limited to small-scale tasks and structurally shallow modifications. As patch size increases or task complexity deepens, performance degrades sharply. Even when a model produces a syntactically valid edit, it often fails to match the structural scope expected of a real engineer. In the next section, we examine whether these structural failures also correspond to deeper semantic errors.
5.4 Semantic Robustness and Failure Characterization
Despite producing syntactically valid Lean code, many model-generated patches fail to fulfill the semantic requirements expressed in the instruction. In this section, we examine the nature of these semantic failures in detail. We analyze patch repairability, categorize common failure modes, and compare error distributions across models to identify systematic weaknesses in current approaches.
Structural Repairability of Model-Generated Patches.
Raw patches generated by LLMs often contain formatting inconsistencies, invalid anchors, or malformed context that prevent successful application. To assess structural robustness, we measure patch application success before and after applying two forms of automated repair: exact match and fuzzy match (See Appendix A for full technical details.). As shown in Table 3, raw apply success rates remain low for all models, ranging from 12% to 16%, indicating poor structural precision in the initial output. Exact match repair offers moderate improvement for DeepSeek models, but fails to resolve structural errors in o3-mini or GPT-4o outputs. Fuzzy match repair significantly boosts success across the board, particularly for DeepSeek R1, Doubao, and Claude Sonnet 3.7. These results highlight the necessity of structural repair mechanisms in enabling meaningful evaluation and practical usability.

### v2

@@ 5.3 Structural Robustness and Task-Scale Sensitivity
To assess the structural robustness of current models, we group benchmark tasks by the size of their ground-truth patches and evaluate pass@16 success rates within each bin. As shown in Figure 6, most models experience a sharp drop in performance as the number of modified lines increases. On small-scale edits (1–5 lines), several models achieve moderate to high success rates. However, performance deteriorates significantly on medium (11–20 lines) and large-scale edits (>20 lines), with all models collapsing to near-zero success beyond 50 lines. Gemini 2.5 Pro Preview consistently performs best in the low to medium complexity range, but fails to generalize to structurally demanding scenarios. These results reveal a fundamental scalability challenge: current LLMs are not robust to increasing structural complexity in proof engineering tasks.
Figure 7: Heatmap of evaluation pass@16 success rates across task types, difficulty levels, and mathematical modules. Column widths represent the proportion of each class in the test set. Color intensity indicates model performance.
We further analyze model performance across task category, difficulty level, and mathematical domain. Figure 7 presents a unified heatmap summarizing pass@16 success rates across these axes.
Task Category.
Models generally handle Bug Fix tasks better than other categories, likely due to their localized nature and relatively narrow scope. Refactor tasks reveal greater variation: Gemini and Claude variants achieve relatively strong performance, benefiting from architectural biases toward structural preservation, while GPT-4o and Doubao consistently underperform. Feature addition is the most challenging category across all models, requiring the introduction of new formal elements while maintaining global coherence.
Difficulty Level.
A steep capability cliff is evident across difficulty levels. Most models perform adequately on Easy tasks, but success rates fall sharply for Medium tasks and approach zero on Hard tasks. These harder tasks typically involve extensive edits, multi-step dependencies, or cross-declaration modifications, indicating a fundamental weakness in current models’ ability to reason over large or interdependent proof contexts.
Mathematical Domain.
Domain-level performance is also highly uneven. Tasks from Dynamics and Tactic modules are consistently easier for all models, likely due to simpler proof structures and shorter edit spans. In contrast, tasks from Combinatorics, Measure Theory, Probability, and Set Theory show uniformly low success rates, reflecting their greater formal density and structural complexity.
Taken together, these results indicate that current LLMs lack structural robustness and task-scale adaptability. While they perform well on localized and syntactically constrained tasks, their effectiveness drops sharply on edits requiring broader semantic planning and structural integration. Overcoming this limitation will require models to incorporate stronger edit planning, context understanding, and global coordination mechanisms, potentially through agentic or feedback-driven architectures.
5.4 Semantic Robustness and Failure Characterization
Error Class Description
Requirement Mismatch Misalignment with the intended instruction
Table 3: Semantic error classes and their typical manifestations.
Figure 8: Distribution of semantic error types across models (log-scaled y-axis). Component omission, quality defects, and incompleteness are the most frequent error types. Claude Sonnet 3.7 (thinking) demonstrates consistently minimal error rates across all categories.
To investigate the gap between syntactic validity and semantic correctness, we conduct a systematic analysis of failure cases using a taxonomy of five high-level semantic error classes, summarized in Table 3. These categories capture recurring patterns where models produce syntactically valid but semantically inadequate patches.
Error Distribution.
Figure 8 illustrates the frequency distribution of semantic error types across different models. Component omission emerges as the most common failure mode, where models neglect to include essential definitions, theorems, or examples required to fulfill the task. This is followed by quality defects, such as redundant or unrefined edits and missing documentation, and incompleteness issues, where edits are only partially implemented or left unfinished.
Model Differences.
GPT-4o and o3-mini exhibit particularly high error rates across all categories, often producing outputs that are not only mathematically incorrect but also misaligned with the original instruction. These models frequently fail to plan edits holistically, leading to missing logic or irrelevant changes. In contrast, Claude Sonnet 3.7 (thinking) shows the lowest error frequency in every category, reflecting strong semantic robustness, structural alignment, and adherence to instructions.
The analysis reveals that semantic failures are not idiosyncratic but follow consistent structural patterns. Most stem from limited reasoning depth, lack of multi-step planning, or failure to reconcile edits with the broader file context. These deficiencies persist even when syntactic correctness is achieved, underscoring the insufficiency of type-checking alone as a proxy for task completion.
5.5 Judge Model Selection and Evaluation Bias
Model Evaluated
Claude
Thinking (%)
Gemini (%)
DeepSeek
R1 (%)
DeepSeek
V3 (%)
Majority
Vote (%)
Gemini 2.5 Pro Preview 18.04 +0.55 16.20 -1.29 17.66 +0.17 18.05 +0.56 17.49
DeepSeek R1 12.55 -0.38 10.56 -2.37 13.80 +0.87 14.80 +1.87 12.93
Claude Sonnet 3.7 (thinking) 12.73 +0.31 11.47 -0.95 12.73 +0.31 12.73 +0.31 12.42
o3-mini 8.60 -2.22 5.94 -4.88 15.31 +4.49 13.42 +2.60 10.82
Claude Sonnet 3.7 10.83 +0.25 8.84 -1.74 11.33 +0.75 11.33 +0.75 10.58
DeepSeek V3 11.81 +0.53 9.44 -1.84 11.93 +0.65 11.92 +0.64 11.28
GPT-4o 6.73 -1.85 5.71 -2.87 13.31 +4.73 8.58 0.00 8.58
Doubao 1.5 Pro 2.98 -0.68 3.48 -0.18 4.08 +0.42 4.08 +0.42 3.66
Squared Deviation from Majority 3.31 78.02 233.61 81.69 -
Family Preference Index (FPI) +0.95% +2.12% –1.16% +1.10% -
Table 4: Judge model comparison: Success rates across different LLM judges. Each judge column shows score and deviation from majority (colored: green for positive, red for negative). The bottom rows report total deviation from consensus (squared) and the Family Preference Index (FPI), which quantifies each judge’s average relative bias toward its own model family.
While syntactic validity can be objectively determined by the Lean compiler, semantic evaluation requires subjective judgment. In our framework, this is delegated to large language models acting as instruction-conditioned judges. However, since these judges are themselves LLMs, a natural concern arises: do they exhibit systematic bias, particularly favoring outputs from models within their own family?
To measure this effect, we introduce a Family Preference Index (FPI), which quantifies the relative deviation between a judge’s ratings of same-family versus other-family models. Formally, for judge $J$,
$\text{FPI}_{J}=\mathbb{E}_{m\in M_{\text{same}}}[\Delta_{J,m}]-\mathbb{E}_{m\in M_{\text{other}}}[\Delta_{J,m}]$
where $\Delta_{J,m}$ is the deviation between $J$’s score and the majority consensus for model $m$. A positive FPI suggests preferential scoring toward same-family outputs.
As summarized in Table 4, most judges display mild but measurable family preferences. Gemini and DeepSeek V3 exhibit positive FPIs, indicating upward bias for their own models. Claude Sonnet 3.7 (thinking) shows the smallest FPI and the lowest deviation from majority voting, reflecting high consistency and relative neutrality. In contrast, DeepSeek R1 reveals a reversed pattern, assigning more favorable ratings to non-family outputs.
While no judge displays extreme favoritism—most FPI values remain within ±2%—such biases can subtly influence evaluations, especially when models are closely matched. These findings motivate our choice of Claude Sonnet 3.7 (thinking) as the primary semantic evaluator, based on its empirical alignment with consensus and minimal systemic bias. To further improve robustness, future work may incorporate multi-judge aggregation, adversarial calibration, or dedicated reference evaluators.
6 Conclusion
This work introduces the paradigm of Automated Proof Engineering (APE) and presents APE-Bench I, the first large-scale benchmark grounded in real-world proof development workflows from Mathlib4. By reframing formal mathematics as instruction-driven code editing and combining Lean compilation with LLM-based semantic evaluation, APE-Bench I offers a rigorous framework for evaluating LLMs in realistic proof engineering tasks. Our empirical results reveal significant performance gaps among state-of-the-art models, particularly under structurally complex scenarios.
A key limitation of APE-Bench I is its focus on single-file edits. Future work will extend this to APE-Bench II and III, which will address multi-file coordination, project-scale verification, and ultimately, autonomous agents capable of planning, editing, and repairing formal libraries. This benchmark series lays the foundation for developing next-generation systems for practical and scalable formal proof maintenance.
Appendix A Extended Background and Related Work
A.1 Formal Proof Engineering: Practice and Challenges
Real-world formal verification projects, including Lean’s Mathlib [9], Isabelle’s Archive of Formal Proofs, and the seL4 microkernel verification project [8], exemplify the scale and complexity of proof engineering. These libraries often contain thousands of interdependent definitions, theorems, and proofs, where modifying a single element may trigger cascading changes across the codebase. Ensuring logical consistency under such conditions demands significant manual labour and deep structural awareness.
While certain tools-such as iCoq for regression-proving [2]-provide partial automation, they are largely heuristic and require intensive human oversight. To date, there remains a lack of automated systems capable of performing high-fidelity edits in complex proof environments.
A.2 Limitations of Existing Benchmarks
The current gap between theorem-proving benchmarks and practical proof engineering extends beyond the aspects highlighted in Table 1. Most benchmarks focus on mathematical reasoning skills in isolation, but lack the context of evolving codebases, dependency management, and continuous integration requirements common in real-world libraries.
Several recent efforts have explored richer proof generation scenarios, such as incorporating intermediate steps, leveraging auxiliary definitions [4, 6, 11, 12, 5], or modifying existing proofs through refactoring [13] and improvement techniques [1]. However, these approaches are still typically used as tools for data augmentation or pipeline enhancement, rather than serving as rigorous evaluation frameworks that capture realistic development workflows.
A.3 Comparison with Software Engineering Benchmarks: The Case of SWE-Bench
The evolution towards more realistic benchmarks is a shared trend across automated reasoning domains. In software engineering, the introduction of benchmarks like SWE-Bench [7] marked a significant step, moving from isolated code generation tasks to evaluating LLMs on resolving real-world GitHub issues within established Python repositories. SWE-Bench’s reliance on existing unit test suites for verifying functional correctness provides an objective and automated measure of success, offering a valuable paradigm.
APE-Bench draws inspiration from this advancement, aiming to create a similarly realistic and challenging environment for automated proof engineering. Like SWE-Bench, it sources tasks from real-world commit histories and focuses on instruction-driven modifications. However, the distinct nature of formal mathematics compared to general software engineering necessitates adaptations in task structure and evaluation. Table 5 outlines the key distinctions.
Aspect SWE-Bench (Software Eng.) APE-Bench (Proof Eng.)
Core Task Resolve functional issues Implement proof modifications
Syntactic Check Compiler/Interpreter Lean Compiler (type checking)
Semantic Verification Unit tests for functional correctness LLM-as-a-judge for semantic alignment with instructions
Objects Observable behavior, I/O correctness Logical validity, structural integrity, adherence to mathematical intent
Table 5: Comparison between SWE-Bench and APE-Bench.
A primary divergence lies in semantic evaluation. While software functionality can often be robustly verified by unit tests capturing desired behaviors, the semantic correctness of a proof modification is more nuanced. A syntactically valid proof (i.e., one that typechecks via the Lean compiler, analogous to code compilation) may still fail to accurately capture the intended mathematical argument, correctly implement a refactoring goal, or properly introduce a new definition as per the natural language instruction. The abstract logical structures and the specific intent behind a proof change are not always fully verifiable by pre-existing automated tests in the same direct manner.
This distinction motivates APE-Bench’s hybrid evaluation strategy: combining compiler-based syntactic verification with an LLM-as-a-Judge (Section 4.2) to assess semantic fulfillment of the natural language instruction. This approach is a pragmatic response to the unique challenges of evaluating proof engineering tasks. Ultimately, APE-Bench aspires to evolve into a standard environment for developing and benchmarking increasingly sophisticated proof engineering agents, mirroring the role SWE-Bench plays in advancing agentic capabilities for software engineering.
Appendix B Technical Details
B.1 Detailed Commit Mining Methodology
Our benchmark construction pipeline incorporates several specialised heuristics to ensure task quality and diversity:
Semantic Filtering Implementation.
To identify commits with recognisable intent, we analyse commit messages using regular expressions that match conventional prefixes like feat:, fix:, or refactor:. We also exclude commits that alter more than 10 files, as these typically represent mechanical changes rather than substantive engineering. Additionally, we filter out commits that only affect non-Mathlib/ directories (e.g., documentation or build scripts).
Content Quality Metrics.
Beyond simple line counting, we implement structural content analysis to identify semantically trivial changes:
  • Whitespace ratio: We exclude diffs where >80% of changes affect only whitespace or empty lines.
  • Comment-only changes: We detect and filter diffs that modify exclusively comment blocks (using Lean’s comment syntax patterns).
  • Import statement changes: Changes affecting only import order or organisation are excluded.
Chunk Segmentation Strategies.
When segmenting modified files, we employ syntax-aware boundaries:
  • For existing files, we use Lean’s declaration structure to identify logical boundaries.
  • For newly added files, we use heuristics based on blank lines and declaration keywords.
  • When expanding large diffs, we maintain declaration integrity by never splitting across def, theorem, or other key structural elements.
These specialised techniques ensure that the resulting benchmark captures meaningful, self-contained, and structurally coherent proof engineering tasks rather than trivial or fragmented edits.
B.2 Instruction Synthesis Methodology
For instruction generation, we employ a carefully engineered prompt that directs Claude Sonnet 3.7 (thinking mode) to analyse a given diff and extract the underlying developer intent. The prompt specifically targets:
  • Intent Extraction: Identifying the core purpose of the edit by analysing both structural changes and semantic modifications.
  • Imperative Formulation: Reframing the extracted intent as a clear, actionable instruction using imperative voice.
  • Technical Precision: Maintaining mathematical and Lean-specific terminology where appropriate.
  • Scoping Accuracy: Ensuring the instruction’s scope corresponds to the actual changes in the diff.
To validate instruction quality, we implement a reverse verification procedure where another LLM instance evaluates the alignment between the generated instruction and the actual diff. Instructions that do not sufficiently specify the target change or contain ambiguities are regenerated until they meet quality thresholds.
B.3 Statistical Validation of Task Labels
Difficulty Labels Beyond Line Counts.
Figure 4(a) shows that while Easy tasks are generally small, Medium and Hard tasks span a broad range of edit sizes, with no strict separation based on line count. This indicates that difficulty labels capture more than superficial metrics-they reflect underlying semantic and structural complexity. This is further confirmed by object-level patterns in Figure 9. Medium and Hard tasks involve more declarations and larger edits per object, particularly for theorems, suggesting greater logical coordination and proof effort. These trends align with empirical results in Section 5, where model performance declines sharply from Easy to Hard, validating the semantic significance of the difficulty labels.
Functional Category Distinctions.
B.4 Eleanstic: Version-Aware Verification Infrastructure
Eleanstic provides a scalable, version-aware verification system for Lean code through a sophisticated snapshot mechanism. Its implementation consists of three key components:
Snapshot Generation.
For each unique commit, Eleanstic first performs a clean lake build and recursively traverses the resulting worktree. Every file is hashed using SHA-256 and stored in a content-addressable storage (CAS) structure with a two-level hash-based directory layout. In parallel, a binary snapshot is created to record the mapping from relative file paths to content hashes and file types (regular or symbolic). This snapshot, encoded using fixed-width binary records, serves as a complete and portable blueprint of the post-build environment, ensuring reproducibility while minimising metadata overhead.
Our implementation runs on a single server with 128 CPU cores and 1TB of RAM. The initial compilation and snapshot generation for all 3734 commits from the Mathlib4 history required approximately 2 days of processing time. Once generated, these snapshots enable efficient verification of candidate patches, with each patch verification typically completing in seconds.
Snapshot Restoration.
At verification time, the appropriate snapshot is loaded and its contents restored into a temporary, isolated worktree that exactly replicates the post-build file structure of the original compilation. Files are fetched from the global CAS and written to disk. This process is stateless and fully parallelisable, enabling fast switching across historical commits without requiring repeated builds. Empirically, Eleanstic restores a full Lean environment in under one second on SSD-backed systems. To support long-term scalability, Eleanstic performs aggressive global deduplication across all snapshots. Across 3734 commits spanning August 2023 to March 2025, it compresses 15.6 TB of build artifacts into just 1.1 TB, yielding a storage reduction of over 93%. This compression enables persistent caching of every historical state required for evaluation, without exhausting disk resources.
Patch Verification.
After restoring the environment, the model-generated patch is applied to the target Lean file, and the resulting content is saved to a temporary file. Verification follows a strict file-level isolation policy: rooted in the restored Mathlib worktree, the Lean compiler compiles only the temporary file without affecting any files in the restored environment. Additionally, any temporary file that attempts to import the target file is discarded to prevent unintended leakage across file boundaries. A patch is considered syntactically valid if it compiles successfully within the corresponding version of Mathlib4 and does not emit warnings.
Appendix C Additional Experimental Results and Analysis
C.1 Model Performance Across pass@k Values
Figure 10: Model performance across pass@k values. Left: Lean Compile pass@$k$ (syntactic validity). Right: LLM Judge pass@$k$ (semantic correctness conditioned on compilation).
Figure 10 illustrates model behaviour across different pass@$k$ thresholds. The left panel reports syntactic success (Lean Compile), while the right panel shows semantic success (LLM-Judge) conditioned on compilation. Gemini 2.5 Pro Preview achieves the highest semantic pass@$k$ across all values of $k$, reflecting consistent success in generating instruction-aligned patches. On the syntactic side, o3-mini leads at higher $k$ values due to more diverse sampling, though its semantic pass@$k$ grows more slowly in comparison. Claude Sonnet 3.7 (thinking) maintains stable but moderate gains in both dimensions. These trends highlight a persistent syntax–semantics gap: models that perform well in compilation do not always satisfy task intent. Sampling diversity improves raw compilation rates but does not necessarily translate to semantic alignment, underscoring the need for instruction-aware generation strategies.
C.2 Patch Length Alignment with Ground-Truth
@@
Figure 11 evaluates whether model-generated patches exhibit scale-aware behaviour by comparing their length to the corresponding human edits. Most models still tend to over-edit small tasks (fewer than 8 lines of substantive changes, excluding blanks and comments), generating unnecessarily long patches. As task size increases, models like Gemini 2.5 Pro Preview and both Claude Sonnet and DeepSeek variants show strong alignment with the ground truth, maintaining near-diagonal trends and low mean squared error (MSE).
In contrast, GPT-4o, o3-mini and Doubao 1.5 Pro significantly under-edit across the board, especially on larger edits, resulting in high patch-length MSE. This misalignment correlates with a higher frequency of quality defects and component omissions.
These results suggest that while strict length matching is not a prerequisite for semantic correctness, extreme deviation—especially persistent over-editing—can lead to incomplete or misaligned edits. Models that strike a balance between scale-awareness and structural flexibility, such as Claude Sonnet, Gemini and DeepSeek V3, tend to perform more robustly in proof engineering tasks.
C.3 DiffRepair: Patch Normalization System
Model outputs are produced in unified diff format, which often contains structurally imprecise or misaligned edits (e.g., incorrect line anchors, out-of-date contexts). To ensure reliable application and verification, we introduce a post-processing phase called DiffRepair, which rewrites imprecise diffs into structurally consistent patches. This mechanism operates in two phases: first, it localises the intended edit position using fuzzy anchor matching; then it reconstructs the patch with preserved structural and contextual integrity.
Intent Localisation.
Given a noisy diff, DiffRepair first identifies where in the original Lean file the intended modification should take place. It uses a multi-granularity fuzzy matching strategy based on contextual anchors surrounding each diff hunk. Anchors are matched using exact or prefix-based line similarity, normalized token-level matching with whitespace tolerance, and structural proximity emphasising declaration-aligned contexts. Matches are scored using a dynamic programming model optimising for structural alignment and modification preservation.
Patch Reconstruction.
Once the target region is localised, DiffRepair reconstructs a clean diff by re-aligning added and deleted lines to structurally valid positions, augmenting missing context to satisfy unified diff constraints, resolving line number offsets and conflicting hunk overlaps, and merging multiple hunks into a minimal, coherent patch that preserves original semantics. This approach ensures that evaluation focuses on semantic correctness rather than superficial formatting issues.